Jane is 6 years older than her younger sister. After 10 years, the sum of their ages will be 50 years. Find their present ages.

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Let the age of Jane’s younger sister is x.

Age of Jane will be x + 6

As per the question, after 10 years the sum of their ages will be 50.

After 10 years,

age of Jane = x + 16

age of her younger sister = x + 10

Therefore,

x + 16 + x +10 = 50

2x + 26 = 50

2x = 24

x = 12

Thus, the present age of Jane’s younger sister is 12 years.

And of Jane’s is 12+6 = 18 years.


Thanks
SAI ADITYA TADALA IVA
2 years ago
X+(X+6)+10 = 50
2X +16 = 50
2X = 34
X = 17
X+6 = 23
Hence Jane's age = 23
& her younger sister's age = 17